### v1

@@ Abstract
This paper studies nonparametric local (over-)identification, in the sense of Chen and Santos (2018), and the associated semiparametric efficiency in modern causal frameworks. We develop a unified approach that begins by translating structural models with latent variables into their induced statistical models of observables and then analyzes local overidentification through conditional moment restrictions. We apply this approach to three leading models: (i) the general treatment model under unconfoundedness, (ii) the negative control model, and (iii) the long-term causal inference model under unobserved confounding. The first design yields a locally just-identified statistical model, implying that all regular asymptotically linear estimators of the treatment effect share the same asymptotic variance, equal to the (trivial) semiparametric efficiency bound. In contrast, the latter two models involve nonparametric endogeneity and are naturally locally overidentified; consequently, some doubly robust orthogonal moment estimators of the average treatment effect are inefficient. Whereas existing work typically imposes strong conditions to restore just-identification before deriving the efficiency bound, we relax such assumptions and characterize the general efficiency bound, along with efficient estimators, in the overidentified models (ii) and (iii).
@@ 1 Introduction
Modern causal models, however, are typically structural rather than purely statistical: they involve unobserved variables—such as potential outcomes, negative controls, or latent confounders—and identification assumptions formulated as restrictions on the joint distribution of these unobserved variables. Our contribution is to make this distinction explicit and to provide a unified approach to studying local (over-)identification in causal inference. We achieve this by translating each structural model into the induced, observationally equivalent statistical model of observables and then analyzing local identification of the observable distribution. This structural-to-statistical translation, combined with the semiparametric efficiency calculation method for general conditional moment restriction models in Ai and Chen (2012), yields a unified framework for determining efficiency bounds and constructing efficient estimators, thereby avoiding the need for case-by-case analyses across different causal designs.
We use this framework to study three representative modern causal inference models: (i) the general treatment model under unconfoundedness (Ai et al., 2021), (ii) the negative control model (Miao et al., 2018), and (iii) the long-term causal inference model under unobserved confounding (Imbens et al., 2024). The first model is without nonparametric endogeneity: under unconfoundedness, treatment assignment is as good as random given observed covariates, so the causal effects are determined by comparing observable conditional distributions. In contrast, the negative control model and long-term causal inference feature nonparametric endogeneity: the treatment effect is identified through solving a nonparametric instrumental variables (NPIV) type inverse problem, in which the nuisance functions appear inside a conditional expectation operator. As shown by Chen and Santos (2018), nonparametric endogeneity typically implies local over-identification, and efficiency considerations become essential in this case.
@@
The remainder of the paper is organized as follows. Section 2 formally introduces the concept of local (over-)identification and its connection to semiparametric efficiency gains. Sections 3, 4, and 5 examine the general treatment model, negative control, and long-term causal inference models, respectively. Section 6 presents an empirical illustration. Section 7 concludes.
@@ 2 Review: Model over-identification and efficiency gains
where cl denotes the closed linear span of a set. By definition, the tangent space $\bar{T}(P)$ is a subset of $L^{2}_{0}(P)$. Whenever this inclusion holds as equality, we say the distribution $P$ is locally just identified by $\mathbf{P}$.
@@
A distribution $P\in\mathbf{P}$ is locally just identified by $\mathbf{P}$ if $\bar{T}(P)=L^{2}_{0}(P)$. Conversely, $P\in\mathbf{P}$ is locally overidentified by $\mathbf{P}$ if $\bar{T}(P)\subsetneqq L^{2}_{0}(P)$.
It is easy to see that global just-identification implies local just-identification, but global over-identification does not imply local over-identification, while local over-identification implies global over-identification. See Chen and Santos (2018) for details.
A parameter of interest is represented as a functional $\mu:\mathbf{P}\to\mathbb{R}^{d_{\mu}}$. Following (Andrews et al., 2025), we refer to $(\mu,\mathbf{P})$ as an econometric model. A parameter $\mu$ is said to be regular at $P$ if there exists $\psi\in L^{2}_{0}(P)$ such that for every path $\theta\mapsto P_{\theta,g}$ with score $g$,
@@ 3 General treatment model under unconfoundedness
where the subscript UC denotes “unconfoundedness.” Any element $P^{*}\in\mathbf{P}^{*}_{\operatorname{UC}}$ of the structural model is a plausible joint distribution of $(\{Y(t)\}_{t\in\mathcal{T}},T)$ such that the unconfoundedness assumption is satisfied.
@@ 4 Negative control model
    Negative–control outcome: $V(1,z)=V(0,z)=V$ for all $z$.
@@
We introduce some notation. Let $e_{1}=(1,0)$ and $e_{0}=(0,1)$. Define the conditional moment function $\rho(h;Y,V,D,Z,X)=(D(Y-h_{1}(V,X)),(1-D)(Y-h_{0}(V,X)))^{\prime}$. Let $\Sigma(Z,X)=\mathbb{E}[\rho\rho^{\prime}|Z,X]$ denote the conditional variance matrix, and write $\Sigma^{-1}(Z,X)$ for its inverse. For $d=0,1$, define $\Gamma_{d}(Z,X)=\mathbb{E}[(h_{d}(V,X)-\mu_{d})\rho^{\prime}\mid Z,X],$ which captures the conditional covariance between the moment functions, and is used for orthogonalization. Finally, let $L(Z,X)$ be the diagonal matrix with diagonal entries p(Z,X)-p(Z,X) and (1p(Z,X))-(1-p(Z,X)), where $p(Z,X)=\mathbb{E}[D\mid Z,X]$.
@@
$\mathbb{EIF}(\mu_{1})$ $=h_{1}(V,X)-\mu_{1}-\Gamma_{1}\rho+(LTH^{-1}T^{*}(e_{1}-\Gamma_{1}L)^{\prime})\Sigma^{-1}\rho,$
@@
$\mathbb{EIF}(\mu_{0})$ $=h_{0}(V,X)-\mu_{0}-\Gamma_{0}\rho+(LTH^{-1}T^{*}(e_{0}-\Gamma_{0}L)^{\prime})\Sigma^{-1}\rho,$
@@ 5 Long-term causal inference via data combination
The first component of the efficiency bound in (5) coincides with the corresponding term in Theorem 7 of Imbens et al. (2024), whereas the second component generally differs. However, when $K$ is bijective, this second term reduces to the expression in Theorem 7 of Imbens et al. (2024). [^5] A formal proof of this statement is given in the proof of Theorem 5.
@@
6 Empirical illustration
As an empirical illustration, we study local overidentification in a nonparametric IV model based on Dahl and Lochner (2012, 2017), which identifies the causal effect of family income on children’s math and reading achievement using changes in the Earned Income Tax Credit (EITC) as an instrument.
The outcome $Y$ is children’s test score gains between adjacent grades, the treatment $T$ is the corresponding change in after-tax family income, the instrument $Z$ is the simulated change in after-tax income driven solely by legislated EITC rule changes (holding household characteristics fixed), and the control $X$ is the family’s pretax income in the prior period.
Identification relies on the fact that changes in EITC generosity generate exogenous shifts in family after-tax income, conditional on pretax income and household demographics. The NPIV moment condition can be written as $\mathbb{E}\big[Y-h(T,X)\big|Z,X\big]=0$, where $h(T,X)$ is the structural function mapping income changes into achievement changes. Intuitively, while $T$ may be endogenous due to unobserved family shocks correlated with both income and achievement, the simulated policy-driven income change $Z$ is exogenous to such shocks, conditional on $X$.
The target parameter of interest is the average structural derivative $\mu=\mathbb{E}\big[\tfrac{\partial}{\partial t}h(T,X)\big]$, which measures the average causal effect of income increments on children’s achievement growth.
We consider two influence-function–based estimators from Chen et al. (2023): the identity-score (IS) estimator, which is inefficient, and the efficient-score (ES) estimator, which attains the semiparametric efficiency bound. In both cases, the influence function for $\mu$ takes the form
$\frac{\partial}{\partial t}h(T,X)-\kappa(X)\,(Y-h(T,X))-\mu$
where the difference between IS and ES lies in the specification of $\kappa$, as detailed in Chen et al. (2023).
The estimation results and the corresponding Hausman test, reported in Table 1, show that the Hausman test does not reject the model at conventional levels, providing a safeguard for the empirical specification. Importantly, such a test would not be feasible (i.e., have nontrivial power) in locally just-identified models. In cases of model misspecification, one can utilize the theory developed in Ai and Chen (2007).
Table 1: Estimates of the Average Structural Derivative and Hausman Test
Inefficient estimate Efficient estimate Hausman test
Estimate 0.0518 -0.0250 2.0830
(SE) (0.2785) (0.2733) [0.1489]
  • Notes: Table reports the estimated average structural derivative $\mu=\mathbb{E}\big[\tfrac{\partial}{\partial t}h(T,X)\big]$. Standard errors are in parentheses. The Hausman test column reports the test statistic, with the corresponding $p$-value in square brackets.
@@ 7 Conclusion
This paper develops a unified framework for analyzing local identification in modern causal inference by explicitly linking structural models to their corresponding observable statistical models. We show that the general treatment model under unconfoundedness is locally just identified. In this case, all regular, asymptotically linear estimators are first-order equivalent, and the semiparametric efficiency bound is degenerate. In contrast, models that are locally overidentified—such as those involving negative controls or long-term causal inference—allow for efficiency improvements and enable nontrivial specification tests.
These results highlight the central role of local overidentification in modern causal analysis. When developing new causal models, it is essential to carry out an explicit model identification analysis. If the model is locally just identified, effort should focus on higher-order properties of regular estimators, since first-order efficiency is degenerate. When feasible, it is preferable to formulate locally overidentified models, which permit the construction of efficient estimators, support specification testing, and yield richer empirical content.

### v2

@@ Abstract
This paper studies nonparametric local (over-)identification and the semiparametric efficiency in modern causal frameworks. We develop a unified approach that begins by translating structural models with latent variables into their induced statistical models of observables and then analyzes local overidentification through conditional moment restrictions. We apply this approach to three popular classes of causal models: (1) the general treatment model under unconfoundedness; (2) the negative control model, and (3) the long-term causal inference model under unobserved confounding. The first model yields a locally just-identified statistical model, implying that all regular asymptotically linear estimators of the treatment effect have the same asymptotic variance, which equals the (trivial) semiparametric efficient variance bound. In contrast, the latter two models involve nonparametric endogeneity and are naturally locally overidentified; consequently, some doubly robust orthogonal moment estimators of the average treatment effect are inefficient. Whereas existing work typically imposes strong conditions to restore local just-identification to justify the efficiency of their doubly robust orthogonal moment estimators, we characterize the semiparametric efficient variance bounds, along with efficient estimators, for the (locally) overidentified models (2) and (3). A small real data application, along with a simulation study, illustrates the semiparametric efficiency gains in model (3).
@@ 1 Introduction
Modern causal models, however, are typically structural rather than purely statistical: they involve unobserved variables—such as potential outcomes, negative controls, or latent confounders—and identification assumptions formulated as restrictions on the joint distribution of these unobserved variables. Our contribution is to make this distinction explicit and to provide a unified approach to studying local (over-)identification in causal models. We achieve this by translating each structural model into an observationally equivalent statistical model of observables and then analyzing local identification of the observable distribution. This structural-to-statistical translation, combined with the semiparametric efficiency calculation method for general sequential conditional moment restriction models in Ai and Chen (2012), yields a unified framework for determining efficiency bounds and constructing efficient estimators, thereby avoiding the need for case-by-case analyses across different causal designs.
We use this framework to study three representative modern causal inference models: (1) the general treatment model under unconfoundedness (Ai et al., 2021), (2) the negative control model (Miao et al., 2018), and (3) the long-term causal inference model under unobserved confounding (Imbens et al., 2024). The first model is without nonparametric endogeneity: under unconfoundedness, treatment assignment is as good as random given observed covariates, so the causal effects are determined by comparing observable conditional distributions. In contrast, the negative control model and the long-term causal inference feature nonparametric endogeneity: the treatment effect is identified through solving a nonparametric instrumental variables (NPIV) type inverse problem, in which the nuisance functions appear inside a conditional expectation operator. As shown by Chen and Santos (2018), nonparametric endogeneity typically implies local over-identification, and efficiency considerations become essential in this case.
@@
The remainder of the paper is organized as follows. Section 2 formally introduces the concept of local (over-)identification and its connection to semiparametric efficiency gains. Sections 3, 4, and 5 examine the general treatment model, negative control, and long-term causal inference models, respectively. Section 6 presents an empirical illustration along with a simulation study. Section 7 concludes.
@@ 2 Review: Model over-identification and efficiency gains
where cl denotes the closed linear span of a set. By definition, the tangent space $\bar{\mathscr{T}}(P)$ is a subset of $L^{2}_{0}(P)$. Whenever this inclusion holds as equality, we say the distribution $P$ is locally just identified by $\mathbf{P}$.
@@
A distribution $P\in\mathbf{P}$ is locally just identified by $\mathbf{P}$ if $\bar{\mathscr{T}}(P)=L^{2}_{0}(P)$. Conversely, $P\in\mathbf{P}$ is locally overidentified by $\mathbf{P}$ if $\bar{\mathscr{T}}(P)\subsetneqq L^{2}_{0}(P)$.
Under global just-identification, $P$ can be approach from any score direction, and hence local just-identification holds given regularity of the model. However, global over-identification does not imply local over-identification, while local over-identification implies global over-identification. See Chen and Santos (2018) for details.
A parameter of interest is represented as a functional $\mu:\mathbf{P}\to\mathbb{R}^{d_{\mu}}$. Following Andrews et al. (2025), we refer to $(\mu,\mathbf{P})$ as an econometric model. A parameter $\mu$ is said to be regular at $P$ if there exists $\psi\in L^{2}_{0}(P)$ such that for every path $\theta\mapsto P_{\theta,g}$ with score $g\in T(P)$,
@@ 3 General treatment model under unconfoundedness
where the subscript UC denotes “unconfoundedness.” Any element $P^{*}\in\mathbf{P}^{*}_{\operatorname{UC}}$ of the structural model is a plausible joint distribution of $(\{Y(t)\}_{t\in\mathcal{T}},T,X)$ such that the unconfoundedness assumption is satisfied.
@@
Our analysis focuses on finite-dimensional causal parameters. For inference on fully nonparametric objects such as the conditional average treatment effect (CATE) function, see, for example, Chang et al. (2015); Lee et al. (2017).
@@ 4 Negative control model
    Negative–control outcome: $V(d,z)=V$ for all $d,z$.
@@
We introduce some notation. Let $e_{1}=(1,0)$ and $e_{0}=(0,1)$. Define the conditional moment function $\rho(h;Y,V,D,Z,X)=(D(Y-h_{1}(V,X)),(1-D)(Y-h_{0}(V,X)))^{\prime}$. Let $\Sigma(Z,X)=\mathbb{E}[\rho\rho^{\prime}|Z,X]$ denote the conditional variance matrix, and write $\Sigma^{-1}(Z,X)$ for its inverse. For $d=0,1$, define $\Gamma_{d}(Z,X)=\mathbb{E}[(h_{d}(V,X)-\mu_{d})\rho^{\prime}\mid Z,X]\Sigma^{-1}(Z,X),$ which captures the conditional covariance between the moment functions, and is used for orthogonalization. Finally, let $L(Z,X)$ be the diagonal matrix with diagonal entries p(Z,X)-p(Z,X) and (1p(Z,X))-(1-p(Z,X)), where $p(Z,X)=\mathbb{E}[D\mid Z,X]$.
@@
$\mathbb{EIF}(\mu_{1})$ $=h_{1}(V,X)-\mu_{1}-\Gamma_{1}\rho-(LTH^{-1}T^{*}(e_{1}-\Gamma_{1}L)^{\prime})^{\prime}\Sigma^{-1}\rho,$
@@
$\mathbb{EIF}(\mu_{0})$ $=h_{0}(V,X)-\mu_{0}-\Gamma_{0}\rho-(LTH^{-1}T^{*}(e_{0}-\Gamma_{0}L)^{\prime})^{\prime}\Sigma^{-1}\rho,$
@@
In the proof of Theorem 3, we formally demonstrate that the efficient influence function reduces to the one in Cui et al. (2024) when $T$ and $T^{*}$ are bijective.
@@ 5 Long-term causal inference via data combination
The first component of the efficiency bound in (5) coincides with the corresponding term in Theorem 7 of Imbens et al. (2024), whereas the second component generally differs. However, when $K$ is bijective, this second term reduces to the corresponding expression in Theorem 7 of Imbens et al. (2024). [^5] A formal proof of this statement is given in the proof of Theorem 5.
@@
6 Numerical studies
6.1 Empirical application: long-term effect of job training
In this section, we apply the efficient estimator developed in Section 5 to study the long-term effect of job training on earnings. We use the Job Corps (JC) dataset as the experimental sample and the Survey of Income and Program Participation (SIPP) as the observational sample.
The JC dataset is obtained from the R package causalweight (Bodory et al., 2025) and originates from the Job Corps National Study, a large-scale randomized evaluation conducted by the US Department of Labor (Schochet et al., 2008). It reports random assignment to the training program, which we take as the treatment $D$, as well as post-treatment log monthly earnings measured one, two, and three years after assignment, which form the short-term outcomes $S_{1},S_{2},S_{3}$. The covariates include years of schooling, age, race, and sex at birth.
The SIPP data are publicly available from the US Census Bureau. [^6] See the Survey of Income and Program Participation, US Census Bureau, https://www.census.gov/programs-surveys/sipp.html. We focus on the 1996 to 2000 panel to ensure close temporal alignment with the JC study. This panel consists of twelve waves (each wave corresponds to a four-month period), with training-related information reported in the topical module of Wave 2. We define the treatment $D$ as participation in any training intended to help train for a new job or improve skills in the current job during the past year. Short-term outcomes $(S_{1},S_{2},S_{3})$ are defined as earnings measured in Waves 2, 5, and 8. The long-term outcome $Y$ is earnings measured in the final wave of the panel. The set of covariates is chosen to match those used in the JC dataset. In the end, the observational sample contains 28,232 untreated and 9,629 treated observations, while the experimental sample contains 3,663 untreated and 5,577 treated observations.
We implement both the estimator of Imbens et al. (2024) and the efficient estimator proposed in Section 5. The sieve spaces for the nuisance function $h$ are constructed using fourth-order spline bases. The probability $\pi$ is estimated using a sieve logit with spline basis, and the conditional covariance matrix is estimated via $K$-nearest neighbors (KNN). The results are reported in Table 1.
Table 1: Estimates of the long-term effect
IKMW Efficient Hausman test
Estimate 0.198 0.181
(SE) (0.036) (0.028) [0.450]
  • Notes: The table reports estimates of the long-term effect. The first two columns correspond to the estimator of Imbens et al. (2024) and our efficient estimator, respectively. Standard errors are reported in parentheses. The Hausman test column reports $p$ values in square brackets.
Both estimators indicate that job training increases long term earnings by nearly 20%, a magnitude that is broadly consistent with existing evidence in the literature (Schochet et al., 2008). The efficient estimator achieves a reduction in the standard error exceeding 20%, reflecting a substantial gain in statistical precision. The Hausman test yields no evidence against the model specification of this empirical study. [^7] In cases of model misspecification, one can utilize the theory developed in Ai and Chen (2007).
6.2 Simulations
We run a simulation study with data-generating processes (DGP) that follows the simulation study in Imbens et al. (2024). We first generate four covariates $\tilde{X}$ from a multivariate normal distribution with mean zero and correlation coefficients calibrated to match those estimated in the empirical application. We also generate latent unobservables $U$ from a multivariate normal distribution with mean zero, variance $0.5$, and zero covariance across components.
Potential intermediate and final outcomes are generated according to the recursive system
$Y(a)$ $=\tau_{y}a+\alpha_{y}^{\top}\tilde{S}_{3}(a)+\beta_{y}^{\top}\tilde{X}+\gamma_{y}^{\top}U+\epsilon_{y},$
$\tilde{S}_{j}(a)$ $=\tau_{j}a+\alpha_{j}\tilde{S}_{j-1}(a)+\beta_{j}^{\top}\tilde{X}+\gamma_{j}^{\top}U+\epsilon_{j},\qquad j\in\{3,2\},$
$\tilde{S}_{1}(a)$ $=\tau_{1}a+\beta_{1}^{\top}\tilde{X}+\gamma_{1}^{\top}U+\epsilon_{1},$
where $\tau_{y}$ and $\tau_{j}$ are scalars and $(\alpha_{y},\beta_{y},\gamma_{y})$ and $(\alpha_{j},\beta_{j},\gamma_{j})$ are scalars, vectors, or matrices of conformable sizes. Following Imbens et al. (2024), we draw the entries of $\tau_{y}$, $(\tau_{j},\alpha_{y},\beta_{y},\gamma_{y})$, and $(\alpha_{j},\beta_{j},\gamma_{j})$ from the uniform distribution on $[0,1]$ and rescale them so that the $\ell_{2}$ norms of the vectors $(\tau_{j},\alpha_{y},\beta_{y},\gamma_{y})$ and the columns of $(\alpha_{j},\beta_{j},\gamma_{j})$ are equal to $0.5$. The disturbances $\epsilon_{j}$ are independent mean zero Gaussian variables with variance $0.5$. The error term $\epsilon_{y}$ has standard deviation $\exp(|\tilde{X}_{1}|)$ to introduce heteroskedasticity.
In the experimental sample, treatment is randomly assigned with probability one half. In the observational sample, treatment follows a logistic specification $\mathbb{P}(D=1\mid\tilde{X},U,G=O)=\left(1+\exp(\kappa_{1}^{\top}\tilde{X}+\kappa_{2}^{\top}U)\right)^{-1},$ where $(\kappa_{1},\kappa_{2})$ are generated using the same sampling and rescaling scheme. This induces persistent confounding through the dependence of treatment on both observed covariates and latent factors.
To introduce nonlinear bridge functions, we apply the elementwise transformation
$\mathrm{sign}(t)\lvert t\rvert^{q},\qquad q\in\{1,2\},$ (8)
to each component of $\tilde{X},\tilde{S}_{1},\tilde{S}_{2},\tilde{S}_{3}$, yielding the observed variables $(X,S_{1},S_{2},S_{3})$. This transformation is invertible and preserves the one to one correspondence between latent and observed variables. When $q=1$, the bridge functions are linear in $(X,S_{1},S_{2},S_{3})$, whereas for $q=2$ they become nonlinear, allowing us to evaluate performance under both linear and nonlinear specifications. We consider $\lambda\in\{0.33,0.50,0.67\}$ as the share of experimental data in the sample.
The nuisance functions are estimated using spline bases with four knots and cubic degree. Ridge regularization is applied to the outcome regression to incorporate modern machine learning techniques, and the conditional variances are estimated via KNN. [^8] Other choices of the tuning parameters deliver similar results. Table 2 reports the simulation results based on 500 Monte Carlo replications. As shown in the table, the efficient estimator achieves a reduction in variance across all designs. The dispersion of both estimators declines with the sample size at the root-$n$ rate, consistent with the asymptotic theory.
1000 2000 4000
$q$ $\lambda$ SD RMSE SD RMSE SD RMSE
IKMW Eff IKMW Eff IKMW Eff IKMW Eff IKMW Eff IKMW Eff
1 .33 .37 .35 .39 .37 .27 .26 .28 .27 .19 .18 .20 .20
.50 .47 .40 .48 .41 .30 .27 .31 .28 .22 .20 .23 .21
.67 .80 .46 .80 .47 .52 .33 .52 .34 .30 .23 .30 .24
2 .33 .37 .34 .39 .37 .27 .25 .29 .28 .19 .18 .21 .20
.50 .47 .38 .48 .39 .32 .26 .33 .29 .23 .20 .25 .22
.67 .84 .44 .84 .46 .55 .31 .56 .34 .33 .22 .33 .24
Table 2: Simulation results.
Simulation results for the estimator in Imbens et al. (2024) (IKMW) and the efficient estimator proposed in Section 5. The number of replications is 500. The parameter $q$ denotes the transformation applied to the variables in (8) to introduce nonlinearity, and $\lambda$ is the share of experimental data in the sample.
@@ 7 Conclusion
This paper develops a unified framework for analyzing local identification in modern causal inference by explicitly linking structural models to their corresponding observable statistical models. We show that the general nonparametric treatment model under unconfoundedness is locally just identified; consequently, all regular, asymptotically linear estimators are first-order equivalent, and the influence functions are all first-order equivalent to the efficient influence function. In contrast, models that are locally overidentified—such as those involving negative controls or long-term causal inference—allow for efficiency improvements and enable nontrivial specification tests.
These results highlight the central role of local overidentification in modern causal analysis. When developing new causal models, it is essential to carry out an explicit model identification analysis. If the model is locally just identified, effort should focus on higher-order properties of regular estimators as they are all first-order equivalent. When feasible, it is preferable to formulate locally overidentified models, which permit the construction of semiparametric efficient estimators, support specification testing with non-trivial powers, and yield richer empirical content.
For future research, the framework could be extended to accommodate dependent data structures and, for example, to investigate event study approaches to causal inference in time series settings (Linton and Whang, 2007; Han et al., 2016).